# LAURA: Enhancing Code Review Generation with Context-Enriched Retrieval-Augmented LLM

### v1

@@ I Introduction
Code review, a process of having peers manually examine source code changes, is critical to software development. Both open source and industrial software projects conduct code review activities to identify defects and ensure software maintainability [bavota2015four, mcintosh2014impact, sadowski2018modern]. With the increasing growth of software scale and complexity, the drawbacks of modern code reviews, e.g., being time-consuming, cannot be overlooked, and many developers face a substantial review workload [zhou2017scalability]. For example, in 2016, a core developer of the Linux kernel raised the argument [JonathanCorbet]: “We are getting more reviewers, but they are coming in slowly and are not anywhere near enough. As a result, the number of unprocessed patches is on the increase.” To alleviate this burden, a substantial amount of studies has focused on automating the code review process, such as recommending the best reviewers [thongtanunam2015should, sulun2019suggesting, asthana2019whodo, chueshev2020expanding, mirsaeedi2020mitigating, chouchen2021whoreview, kong2022recommending, pandya2022corms], predicting code locations that need review [hellendoorn2021towards], suggesting relevant code review comments [hong2022commentfinder, shuvo2023recommending], and optimizing code before submission [tufano2021towards, tufano2022using].
@@
Driven by the rapid development of artificial intelligence technologies, many studies have shifted towards automatically generating code reviews by leveraging neural machine translation and pre-trained models [tufano2022using, li2022automating, lu2023llama]. However, achieving satisfactory results remains challenging because of the complexity and knowledge-intensive nature of code review. The rise of large language models (LLMs) brings new opportunities for improvement, as they demonstrate exceptional performance in natural language understanding, generation, and a wide range of SE automation tasks [eliseeva2023commit, sun2024gptscan, guo2024exploring]. However, directly using LLMs for code review generation faces several limitations. On one hand, LLMs often lack access to essential contextual information such as pull request (PR) details, code change specifics, and the domain-specific review experience that human reviewers rely on – all of which are crucial for producing high-quality reviews [kononenko2016code, turzo2024makes]. On the other hand, LLMs lack systematic guidance when generating code reviews; they typically cannot understand well the review’s scope and focus, nor do they operate within a well-defined logical framework to structure their feedback. This absence of context and structured guidance can lead to suboptimal review quality.
@@
To overcome these limitations, we propose LAURA, a novel LLM AUgmentation Framework for Code Review GenerAtion. LAURA consists of three components – context augmentation, review exemplar retrieval, and systematic guidance – combining augmentation and prompts to improve LLMs in code review generation. We leverage relevant contextual information of PRs and code changes and construct a history code review dataset to provide LLMs with relevant background knowledge and review experience. Specifically, since existing code review datasets do not have the context information we require and tend to have a quarter of low-quality data [tufano2024code], we construct a new dataset with careful filtering. Our dataset contains 301,256 diff-comment-info series, which includes complete PR data from 1,807 popular GitHub projects. We manually annotate 384 recent entries for evaluation, while the remaining 298,494 entries dated before December 26, 2024, are used as the retrieval-augmented generation (RAG) database to provide valuable reference reviews. This enables the LLM to leverage relevant context and review experience, similar to a human reviewer. To address the second limitation – lack of guidance – we design prompts that provide clear instructions and a logical review structure, helping LLMs better understand review focus and generate coherent, high-quality comments.
@@
We apply the LAURA framework to two well-performed LLMs, i.e., ChatGPT-4o and DeepSeek v3 (hereinafter, we name them as LAURA-GPT and LAURA-DS, respectively), and conduct experiments to validate the effectiveness of LAURA. To overcome the limitations of syntactic similarity-based metrics in evaluating LLM-generated text [fiske1992but], we conduct both LLM-based and human evaluations. The results show that LAURA-GPT and LAURA-DS consistently rank at the top across four out of five LLM evaluation metrics for assessing the overall review generation performance – readability, relevance, sufficiency, and operability – only slightly trailing in brevity. Additionally, both models achieve I-Score and IH-Score values of 20.0%, 42.2% and 18.5%, 40.4% respectively, in human evaluation metrics, which measure the usefulness of the generated content, meaning that both models effectively identified issues in about one-fifth of cases and provided at least helpful review comments in over two-fifths of cases. Compared to their respective base models, LAURA-GPT and LAURA-DS improve their I-Score by 30.5% and 36.6%, and their IH-Score by 37.3% and 38.4%, respectively. This indicates a significant improvement over directly using ChatGPT-4o and DeepSeek v3. LAURA also significantly outperforms CodeReviewer [li2022automating], the pre-trained SOTA approach in code review generation. Furthermore, we conduct an ablation study to analyze the impact of LAURA’s three components on code review quality, confirming the effectiveness of the approach.
@@
  • We have made the retrieval-augmented data publicly available to facilitate future improvements in code review generation [LAURAfigshare].
@@ II-A Context Augmentation
Commit message. Each commit within a PR may cover different aspects of the changes, with messages documenting those changes [tian2022what]. Commit messages typically contain less information but are more fine-grained. Including the commit message can provide LLMs with more commit-specific contextual information and help them better understand the changes.
Changed file path. Similarly, the file path provides more fine-grained information compared to the commit message, hinting at the file’s role and the nature of the changes. Prior work shows paths can aid in review tasks [shuvo2023recommending]. It could help LLMs assess change types and locate issues more effectively.
@@
For each code change, we gather the relevant information mentioned earlier and provide all of it, except for code context, directly to the LLMs. We integrate the code context into the diff as follows: (1) We use Tree-sitter [Tree-sitter] to build an Abstract Syntax Tree (AST) of the source file corresponding to the diff. If possible, we extend the diff to the boundaries of the enclosing function/method, limited to three times the original diff length; otherwise, we extend it up to that length. (2) We extract all header/library/package imports appearing before the diff and include any not already in the expanded diff. (3) We annotate each line of the extended diff with line numbers: added lines as “[line number]+”, unchanged lines as “[line number] ”, and deleted lines as “[]-”. In rare cases (under 3%) where the source file retrieved via the GitHub GraphQL API doesn’t match the diff exactly, we only apply step (3). This approach enables us to create context-extended diffs, giving LLMs richer and more useful code context information.
@@ II-B Review Exemplar Retrieval
Existing studies indicate that the types of issues identified during code reviews and the level of usefulness of the reviews are related to the experience and knowledge of reviewers [kononenko2016code, turzo2024makes]. Simply relying on LLMs can hardly grasp the pertinent knowledge from the vast amount of pre-training data. Thus, we hypothesize that providing LLMs with diffs similar to the code diff under review, along with the review comments made by other reviewers for those diffs, can serve as a reference for generating review comments. Prior studies [wang2021context, zhang2024rag] applied retrieval-augmented approaches to automatically generate commit messages and achieved promising effectiveness. In this study, we first introduce the retrieval-augmented generation (RAG) method to retrieve diffs similar to the code diff under review and their corresponding review comments, using this retrieved information as part of the input to LLMs to achieve better quality in the generated review comments.
We use cosine similarity scoring [ochiai1957zoogeographic], based on transformer encoders and vector embeddings, to assess code diff similarity and retrieve the most relevant code diffs and their review comments. For embedding, we use CodeT5+ [wang2023codet5plus], an enhanced version of CodeT5 [wang2021codet5], known for its Transformer encoder-decoder architecture and rich code representations from pre-training. We specifically use the “codet5p-110m-embedding” model to convert code diffs into 256-dimensional dense vectors, which encapsulate semantic information. Kartal et al. [kartal2024automating] show that transformers, when pre-trained and fine-tuned, outperform traditional methods like Word2Vec and Code2Vec in encoding embeddings, making them well-suited for retrieval-augmented generation (RAG). After embedding the code diffs, we compute cosine similarity scores to identify the most relevant code diff and its review comments, which will be used to enhance the performance of LLMs.
@@ II-C Systematic Guidance
We provide the augmented information, i.e., PR context and retrieved review exemplars, to LLMs along with the code diff under review through prompts. However, we still need to address the challenge that LLMs lack scientific guidance. Research shows that a well-designed prompt can guide LLMs in understanding the task and focusing on key points, leading to more structured, detailed, and accurate outputs, with its importance potentially surpassing that of the data itself [white2023prompt]. Many automated tasks in software engineering have proved this, such as [bae2024enhancing, ahmed2024automatic, pornprasit2024fine]. Thus, we also explore systematic guidance to enhance the performance of LLMs in the code review generation task. Inspired by relevant studies [white2023prompt, amatriain2024prompt], we design three prompt components – diff example, review guidelines, and chain of thought – to improve the generation quality of LLMs, focusing on input comprehension, task-critical point emphasis, and structured logical guidance, respectively.
@@
Chain of thought. LLMs also need guidance on reasoning and information use. Prior work shows the chain of thought (CoT) approach can improve review outcomes [tufano2024code]. Thus, we adopt CoT to provide LLMs with a clear logical framework, guiding them to logically process the augmented information: first, understanding the intent and content of changes, then identifying and explaining issues, and finally offering suggestions. We also integrate contextual information into the CoT to help LLMs better leverage relevant details during review.
@@ III-A Research Questions
RQ1: How does LAURA perform on code review generation compared to the baselines? In this RQ, we evaluate our method using ChatGPT-4o and DeepSeek v3 to generate review comments on our evaluation dataset, as they represent closed-source and open-source LLMs, respectively. We compare two setups: (1) LAURA guidance, and (2) direct generation, and benchmark against CodeReviewer [li2022automating]. Results are assessed via LLM and human evaluation.
@@ III-B Dataset Construction
Only a few code review generation datasets are available, notably the T5CR dataset [tufano2022using] and the CodeReviewer dataset [li2022automating]. Many studies rely on the CodeReviewer dataset, but existing datasets suffer from quality issues. Tufano et al. found that around 25% of these datasets are of low quality, with the CodeReviewer dataset [tufano2024code] reaching about 32%. This significantly limits dataset reuse. Furthermore, these datasets lack contextual information beyond the code itself and remove the source links of PRs, further constraining their usefulness. Therefore, we construct our own code review dataset with careful filtering to ensure high quality.
@@ III-B2 Data Collection
Table I summarizes our data collection and cleaning process. We primarily used the GitHub GraphQL API [GitHubGraphQL] to retrieve pull request data, as its flexible query system allows precise JSON data retrieval via HTTP POST requests. We collected PR titles, bodies, numbers, states, authors, reviewers, review comments (with timestamps), commit messages, diffs, and file paths – collectively referred to as a “diff-comment-info series.” We filtered out cases where the PR author and reviewer were the same. Since comment-associated diffs were truncated, we fetched complete diffs via the GitHub commit-comparing function [GitHubCompare].
To improve data quality, we applied the 10-line rule [bosu2015characteristics], which assumes that code changes within 10 lines of a review location likely address the review and thus indicate valuable feedback. Following Rong et al. [rong2024distilling], we sorted commits and reviews chronologically within PRs and checked whether later commits modified code near earlier review locations. We kept only data conforming to the 10-line rule. After initial processing, we obtained 1,020,537 diff-comment-info series, but further filtering was necessary as the 10-line rule alone does not guarantee high quality.
@@ III-B3 Data Filtering
After data collection, we cleaned the collected data. First, we applied rule-based filtering: (1) We removed bot-generated comments by filtering reviewer names with “bot”, “-cr”, or “gpt” suffixes, and by using Golzadeh et al.’s bot list [golzadeh2021ground]. (2) We filtered invalid and irrelevant comments by (a) removing comments with two or fewer words, as they typically offer little value [dyer2013boa, schall2024commitbench], and (b) applying Tufano et al.’s heuristic algorithm based on manually designed patterns [tufano2022using]. (3) We removed comments containing non-ASCII characters. After these steps, we retained 817,833 diff-comment-info series.
Next, we applied an LLM-based filter to improve review quality. As Tufano et al. [tufano2024code] note, 25% of existing review comments remain low-quality despite basic filtering. We used GPT-4o mini to evaluate each comment, its corresponding diff, and a task prompt: “Determine whether the review comments provided are valuable for improving the code diff. If the review comments point out problems in the code diff or give useful improvement suggestions, and are easy to understand, answer ‘yes’; if the review comments are irrelevant to the content of the code diff (such as only requiring it to be done in a separate PR or commit), or are difficult to understand or not clearly described, answer ‘no’. The answer should be one of ‘yes’ or ‘no’, without other content.”
@@ III-B4 Retrieval System Construction
We split the constructed dataset by time and selected 298,494 diff-comment-info samples prior to December 26, 2024, as the retrieval-augmented generation database. To improve retrieval efficiency, we pre-embedded the code diffs for similarity comparison using the “codet5p-110m-embedding” model [wang2023codet5plus], generating 256-dimensional vectors. For test diffs, we embedded them with the same model and calculated cosine similarity scores against the RAG database. The most similar diffs and their review comments were merged into the test diff-comment-info series as review references for the LLM. With data partitioning and post-processing complete, we finalized the review exemplar retrieval component for our composite method.
@@ III-B5 Evaluation Dataset Construction
Since LAURA involves LLM augmentation and human evaluation, large-scale experiments aren’t feasible. We intended to sample 384 instances for evaluation (95% confidence level, 5% margin of error). Given the possibility of residual low-quality data in the LLM-filtered set and the risk of bias from discarding valuable comments, we opted to manually annotate high-quality data from the dataset without LLM filtering. We manually annotated high-quality code review comments based on the following criteria: (1) Readability and comprehensibility (this criterion excluded comments like “s/Check less/Check if less/” that were difficult to understand); (2) Clear relevance to issues in the code under review (this criterion excluded comments with low relevance to the code itself, such as “I don’t think this should be part of this PR?”); (3) Containing at least an effective issue explanation or code suggestion (this criterion excluded comments with low informational value, such as “When does this happen?” or “Thanks”); (4) Self-contained understanding (this criterion excluded comments like “Ditto, also change this” that lacked essential contextual information). The annotation continues until we obtain 384 valid samples. To minimize data leakage risk (i.e., the chance that LLMs had seen the data during training), we only selected the diff-comment-info series dated after December 26, 2024 (the release date of DeepSeek v3, the latest model used in our experiments). We merged and shuffled the data. Then, the first two authors independently annotated each sample. A Cohen’s kappa [cohen1960coefficient] of 0.883 between their results, indicating a high level of inter-rater agreement. Subsequently, any disagreements were resolved by conducting several meetings and involving a third author as arbiter. In total, we annotated 546 samples, yielding 384 high-quality samples for evaluation.

### v2

@@ I Introduction
Code review, a process of having peers manually examine source code changes, is critical to software development. Both open source and industrial software projects conduct code review activities to identify defects and ensure software maintainability [5, 34, 43]. With the increasing growth of software scale and complexity, the drawbacks of modern code reviews, e.g., being time-consuming, cannot be overlooked, and many developers face a substantial review workload [66]. For example, in 2016, a core developer of the Linux kernel raised the argument [10]: “We are getting more reviewers, but they are coming in slowly and are not anywhere near enough. As a result, the number of unprocessed patches is on the increase.” To alleviate this burden, a substantial amount of studies has focused on automating the code review process, such as recommending the best reviewers [49, 47, 3, 8, 35, 7, 23, 38], predicting code locations that need review [19], suggesting relevant code review comments [20, 46], and optimizing code before submission [54, 53].
@@
Driven by the rapid development of artificial intelligence technologies, many studies have shifted towards automatically generating code reviews by leveraging neural machine translation and pre-trained models [53, 27, 32]. However, achieving satisfactory results remains challenging because of the complexity and knowledge-intensive nature of code review. The rise of large language models (LLMs) brings new opportunities for improvement, as they demonstrate exceptional performance in natural language understanding, generation, and a wide range of SE automation tasks [12, 48, 18]. However, directly using LLMs for code review generation faces several limitations. On one hand, LLMs often lack access to essential contextual information such as pull request (PR) details, code change specifics, and the domain-specific review experience that human reviewers rely on – all of which are crucial for producing high-quality reviews [24, 55]. On the other hand, LLMs lack systematic guidance when generating code reviews; they typically cannot understand well the review’s scope and focus, nor do they operate within a well-defined logical framework to structure their feedback. This absence of context and structured guidance can lead to suboptimal review quality.
@@
To overcome these limitations, we propose LAURA, a novel LLM AUgmentation Framework for Code Review GenerAtion. LAURA consists of three components – context augmentation, review exemplar retrieval, and systematic guidance – combining augmentation and prompts to improve LLMs in code review generation. We leverage relevant contextual information of PRs and code changes and construct a history code review dataset to provide LLMs with relevant background knowledge and review experience. Specifically, since existing code review datasets do not have the context information we require and tend to have a quarter of low-quality data [52], we construct a new dataset with careful filtering. Our dataset contains 301,256 diff-comment-info series, which includes complete PR data from 1,807 popular GitHub projects. We manually annotate 384 recent entries for evaluation, while the remaining 298,494 entries dated before December 26, 2024, are used as the retrieval-augmented generation (RAG) database to provide valuable reference reviews. This enables the LLM to leverage relevant context and review experience, similar to a human reviewer. To address the second limitation – lack of guidance – we design prompts that provide clear instructions and a logical review structure, helping LLMs better understand review focus and generate coherent, high-quality comments.
@@
We apply the LAURA framework to two well-performed LLMs, i.e., ChatGPT-4o and DeepSeek v3 (hereinafter, we name them as LAURA-GPT and LAURA-DS, respectively), and conduct experiments to validate the effectiveness of LAURA. To overcome the limitations of syntactic similarity-based metrics in evaluating LLM-generated text [13], we conduct both LLM-based and human evaluations. The results show that LAURA-GPT and LAURA-DS consistently rank at the top across four out of five LLM evaluation metrics for assessing the overall review generation performance – readability, relevance, sufficiency, and operability – only slightly trailing in brevity. Additionally, both models achieve I-Score and IH-Score values of 20.0%, 42.2% and 18.5%, 40.4% respectively, in human evaluation metrics, which measure the usefulness of the generated content, meaning that both models effectively identified issues in about one-fifth of cases and provided at least helpful review comments in over two-fifths of cases. Compared to their respective base models, LAURA-GPT and LAURA-DS improve their I-Score by 30.5% and 36.6%, and their IH-Score by 37.3% and 38.4%, respectively. This indicates a significant improvement over directly using ChatGPT-4o and DeepSeek v3. LAURA also significantly outperforms CodeReviewer [27], the pre-trained SOTA approach in code review generation. Furthermore, we conduct an ablation study to analyze the impact of LAURA’s three components on code review quality, confirming the effectiveness of the approach.
@@
  • We have made the retrieval-augmented data publicly available to facilitate future improvements in code review generation [64].
@@ II-A Context Augmentation
Commit message. Each commit within a PR may cover different aspects of the changes, with messages documenting those changes [50]. Commit messages typically contain less information but are more fine-grained. Including the commit message can provide LLMs with more commit-specific contextual information and help them better understand the changes.
Changed file path. Similarly, the file path provides more fine-grained information compared to the commit message, hinting at the file’s role and the nature of the changes. Prior work shows paths can aid in review tasks [46]. It could help LLMs assess change types and locate issues more effectively.
@@
For each code change, we gather the relevant information mentioned earlier and provide all of it, except for code context, directly to the LLMs. We integrate the code context into the diff as follows: (1) We use Tree-sitter [51] to build an Abstract Syntax Tree (AST) of the source file corresponding to the diff. If possible, we extend the diff to the boundaries of the enclosing function/method, limited to three times the original diff length; otherwise, we extend it up to that length. (2) We extract all header/library/package imports appearing before the diff and include any not already in the expanded diff. (3) We annotate each line of the extended diff with line numbers: added lines as “[line number]+”, unchanged lines as “[line number] ”, and deleted lines as “[]-”. In rare cases (under 3%) where the source file retrieved via the GitHub GraphQL API doesn’t match the diff exactly, we only apply step (3). This approach enables us to create context-extended diffs, giving LLMs richer and more useful code context information.
@@ II-B Review Exemplar Retrieval
Existing studies indicate that the types of issues identified during code reviews and the level of usefulness of the reviews are related to the experience and knowledge of reviewers [24, 55]. Simply relying on LLMs can hardly grasp the pertinent knowledge from the vast amount of pre-training data. Thus, we hypothesize that providing LLMs with diffs similar to the code diff under review, along with the review comments made by other reviewers for those diffs, can serve as a reference for generating review comments. Prior studies [56, 63] applied retrieval-augmented approaches to automatically generate commit messages and achieved promising effectiveness. In this study, we first introduce the retrieval-augmented generation (RAG) method to retrieve diffs similar to the code diff under review and their corresponding review comments, using this retrieved information as part of the input to LLMs to achieve better quality in the generated review comments.
We use cosine similarity scoring [36], based on transformer encoders and vector embeddings, to assess code diff similarity and retrieve the most relevant code diffs and their review comments. For embedding, we use CodeT5+ [57], an enhanced version of CodeT5 [58], known for its Transformer encoder-decoder architecture and rich code representations from pre-training. We specifically use the “codet5p-110m-embedding” model to convert code diffs into 256-dimensional dense vectors, which encapsulate semantic information. Kartal et al. [22] show that transformers, when pre-trained and fine-tuned, outperform traditional methods like Word2Vec and Code2Vec in encoding embeddings, making them well-suited for retrieval-augmented generation (RAG). After embedding the code diffs, we compute cosine similarity scores to identify the most relevant code diff and its review comments, which will be used to enhance the performance of LLMs.
@@ II-C Systematic Guidance
We provide the augmented information, i.e., PR context and retrieved review exemplars, to LLMs along with the code diff under review through prompts. However, we still need to address the challenge that LLMs lack scientific guidance. Research shows that a well-designed prompt can guide LLMs in understanding the task and focusing on key points, leading to more structured, detailed, and accurate outputs, with its importance potentially surpassing that of the data itself [60]. Many automated tasks in software engineering have proved this, such as [4, 1, 39]. Thus, we also explore systematic guidance to enhance the performance of LLMs in the code review generation task. Inspired by relevant studies [60, 2], we design three prompt components – diff example, review guidelines, and chain of thought – to improve the generation quality of LLMs, focusing on input comprehension, task-critical point emphasis, and structured logical guidance, respectively.
@@
Chain of thought. LLMs also need guidance on reasoning and information use. Prior work shows the chain of thought (CoT) approach can improve review outcomes [52]. Thus, we adopt CoT to provide LLMs with a clear logical framework, guiding them to logically process the augmented information: first, understanding the intent and content of changes, then identifying and explaining issues, and finally offering suggestions. We also integrate contextual information into the CoT to help LLMs better leverage relevant details during review.
@@ III-A Research Questions
RQ1: How does LAURA perform on code review generation compared to the baselines? In this RQ, we evaluate our method using ChatGPT-4o and DeepSeek v3 to generate review comments on our evaluation dataset, as they represent closed-source and open-source LLMs, respectively. We compare two setups: (1) LAURA guidance, and (2) direct generation, and benchmark against CodeReviewer [27]. Results are assessed via LLM and human evaluation.
@@ III-B Dataset Construction
Only a few code review generation datasets are available, notably the T5CR dataset [53] and the CodeReviewer dataset [27]. Many studies rely on the CodeReviewer dataset, but existing datasets suffer from quality issues. Tufano et al. found that around 25% of these datasets are of low quality, with the CodeReviewer dataset [52] reaching about 32%. This significantly limits dataset reuse. Furthermore, these datasets lack contextual information beyond the code itself and remove the source links of PRs, further constraining their usefulness. Therefore, we construct our own code review dataset with careful filtering to ensure high quality.
@@ III-B2 Data Collection
Table I summarizes our data collection and cleaning process. We primarily used the GitHub GraphQL API [16] to retrieve pull request data, as its flexible query system allows precise JSON data retrieval via HTTP POST requests. We collected PR titles, bodies, numbers, states, authors, reviewers, review comments (with timestamps), commit messages, diffs, and file paths – collectively referred to as a “diff-comment-info series.” We filtered out cases where the PR author and reviewer were the same. Since comment-associated diffs were truncated, we fetched complete diffs via the GitHub commit-comparing function [15].
To improve data quality, we applied the 10-line rule [6], which assumes that code changes within 10 lines of a review location likely address the review and thus indicate valuable feedback. Following Rong et al. [42], we sorted commits and reviews chronologically within PRs and checked whether later commits modified code near earlier review locations. We kept only data conforming to the 10-line rule. After initial processing, we obtained 1,020,537 diff-comment-info series, but further filtering was necessary as the 10-line rule alone does not guarantee high quality.
@@ III-B3 Data Filtering
After data collection, we cleaned the collected data. First, we applied rule-based filtering: (1) We removed bot-generated comments by filtering reviewer names with “bot”, “-cr”, or “gpt” suffixes, and by using Golzadeh et al.’s bot list [17]. (2) We filtered invalid and irrelevant comments by (a) removing comments with two or fewer words, as they typically offer little value [11, 45], and (b) applying Tufano et al.’s heuristic algorithm based on manually designed patterns [53]. (3) We removed comments containing non-ASCII characters. After these steps, we retained 817,833 diff-comment-info series.
Next, we applied an LLM-based filter to improve review quality. As Tufano et al. [52] note, 25% of existing review comments remain low-quality despite basic filtering. We used GPT-4o mini to evaluate each comment, its corresponding diff, and a task prompt: “Determine whether the review comments provided are valuable for improving the code diff. If the review comments point out problems in the code diff or give useful improvement suggestions, and are easy to understand, answer ‘yes’; if the review comments are irrelevant to the content of the code diff (such as only requiring it to be done in a separate PR or commit), or are difficult to understand or not clearly described, answer ‘no’. The answer should be one of ‘yes’ or ‘no’, without other content.”
@@ III-B4 Retrieval System Construction
We split the constructed dataset by time and selected 298,494 diff-comment-info samples prior to December 26, 2024, as the retrieval-augmented generation database. To improve retrieval efficiency, we pre-embedded the code diffs for similarity comparison using the “codet5p-110m-embedding” model [57], generating 256-dimensional vectors. For test diffs, we embedded them with the same model and calculated cosine similarity scores against the RAG database. The most similar diffs and their review comments were merged into the test diff-comment-info series as review references for the LLM. With data partitioning and post-processing complete, we finalized the review exemplar retrieval component for our composite method.
@@ III-B5 Evaluation Dataset Construction
Since LAURA involves LLM augmentation and human evaluation, large-scale experiments aren’t feasible. We intended to sample 384 instances for evaluation (95% confidence level, 5% margin of error). Given the possibility of residual low-quality data in the LLM-filtered set and the risk of bias from discarding valuable comments, we opted to manually annotate high-quality data from the dataset without LLM filtering. We manually annotated high-quality code review comments based on the following criteria: (1) Readability and comprehensibility (this criterion excluded comments like “s/Check less/Check if less/” that were difficult to understand); (2) Clear relevance to issues in the code under review (this criterion excluded comments with low relevance to the code itself, such as “I don’t think this should be part of this PR?”); (3) Containing at least an effective issue explanation or code suggestion (this criterion excluded comments with low informational value, such as “When does this happen?” or “Thanks”); (4) Self-contained understanding (this criterion excluded comments like “Ditto, also change this” that lacked essential contextual information). The annotation continues until we obtain 384 valid samples. To minimize data leakage risk (i.e., the chance that LLMs had seen the data during training), we only selected the diff-comment-info series dated after December 26, 2024 (the release date of DeepSeek v3, the latest model used in our experiments). We merged and shuffled the data. Then, the first two authors independently annotated each sample. A Cohen’s kappa [9] of 0.883 between their results, indicating a high level of inter-rater agreement. Subsequently, any disagreements were resolved by conducting several meetings and involving a third author as arbiter. In total, we annotated 546 samples, yielding 384 high-quality samples for evaluation.